Find a homogeneous linear differential equation with constant coefficient that is satisfied by y=4xe^xsin2x
1
Expert's answer
2020-12-07T20:08:07-0500

"\displaystyle
y = 4xe^x\sin(2x)\\

\textsf{If}\,\, y = e^{\alpha x}(A\cos(\beta x) + B\sin(\beta x)),\\
\textsf{the auxiliary equation is}\\

m = \alpha \pm j\beta\\

\therefore m = 1 \pm j2\\

\textsf{The quadratic equation whose roots are}\,\,
1 \pm j2 \,\, \textsf{are}\\

m^2 - (1 + j2 + 1 - j2)m + (1 + j2)(1 - j2) = 0\\

m^2 - 2m + 5 = 0\\

\textsf{For}\,\, e^x\sin(2x) \,\, \textsf{to be multiplied by}\,\, x, \\
\textsf{the roots of}\,\,m\,\,\textsf{must be repeated.}\\

\therefore m = 1 \pm j2 \,\,\textsf{twice}\\

(m^2 - 2m + 5)^2 = 0\\

m^4 - 4m^3 + 14m^2 - 20m + 25 = 0\\

\therefore \textsf{The homogenous differential equation is}\\
y^{(iv)} - 4y^{(iii)} + 14y'' - 20y' + 25y = 0"